Can somebody help me with this question? I need it answer by today.
A financial manager invested 20% of a client's money in bonds paying 9% annual simple interest, 35% in an 8% simple interest account, and the remainder in 9.5% corporate bonds. Find the amount invested in each if the total annual interest earned is $4,970.
at 9% $
at 8% $
at 9.5% $

Answer by Alan3354(69443)

A financial manager invested 20% of a client's money in bonds paying 9% annual simple interest, 35% in an 8% simple interest account, and the remainder in 9.5% corporate bonds.
Find the amount invested in each if the total annual interest earned is $4,970.
The remainder = 45%
A = total amount
-----------------
0.2*A*0.09 + 0.35*A*0.08 + 0.45*A*0.095 = 4970
A*(0.018 + 0.028 + 0.04275) = A*0.08875 = 4970
A = 4970/0.08875
A = $56000